OpenStudy (anonymous):

Simplify: (5.4 x 103)(7.2 x 10-6) 3.888 x 1017 3.888 x 10-3 3.888 x 10-2 3.888 x 1010

OpenStudy (diyadiya):

First Multiply 5.4 x 7.2

OpenStudy (diyadiya):

Then multiply \( \large 10^3 \times 10^{-6} \)

OpenStudy (anonymous):

is it the last one

OpenStudy (diyadiya):

for that use this formula \[\Large x^m \times x^n = x^{m+n}\] for example \[\Large 10^2 \times 10^7 = 10^{2+7} =10^9\]

OpenStudy (anonymous):

so im i right

OpenStudy (diyadiya):

How did you get that?

OpenStudy (anonymous):

I simplify it is the answer wrong

OpenStudy (diyadiya):

Last one is not the answer if you show me your work i can tell you where you went wrong